OpenStudy (zzr0ck3r):

Consider an animal population P(t) with constant death rate s = 0.02 ( deaths per animal per month) and with birth rate B proportional to P. Suppose that P(0) = 200 and P`(0) = 2. When is P = 1000?

OpenStudy (paxpolaris):

\[\large P'(t)=B(t)-s\] and, \(\large \because B(t) \propto P(t) \implies B(t) = k \cdot P(t) \) for some constant k

OpenStudy (paxpolaris):

solve for k using \[\large P'(0)= k \cdot P(0)-0.02\]

OpenStudy (zzr0ck3r):

what is the B(t) alpha P(t)?

OpenStudy (paxpolaris):

rate of change of population = birth rate - death rate you are given that birth rate “B(t)” is proportional to P(t)

OpenStudy (zzr0ck3r):

ahh alpha is praportional I see

OpenStudy (zzr0ck3r):

tyvm

OpenStudy (paxpolaris):

i'm getting \(\large k=0.0101\) i hope u can take it from here.. i'm a bit rusty on differential equations

OpenStudy (zzr0ck3r):

yeah for susre, just wanted to know what it was saying:)